Alicia recieved some money from work. She spent one-fifth of it on gas. Then she spent half of what she had left on clothes. She spent seven dollars at lunch. When she got home she had thirteen dollars left. How much did she have to start with?
Answer by checkley75(3666)

4/5x(.5)-7=13
2x/5=20
2x=100
x=100/2
x=50 amount of money she started with.
proof
4/5*50(.5)-7=13
200/5*.5=13+7
40*.5=20
20=20